# Band inversion transition in HgTe nanowire grown along the [001] direction

### v1

@@ IV Transformation of representation
We find $\Delta^{\prime}_{e}$ is always larger than $\Delta^{\prime}_{h1}$, regardless of the detailed values of $\Delta_{e,h_{1}}$ and $\zeta_{e,h_{1},eh_{1}}$. This result has actually been manifested in Fig. 1(b), where the $E_{1}$ subband always lies above the $H_{1}$ subband at $k_{z}\!=\!0$ when they are plotted as functions of the nanowire radius $R$. Hence, the anticrossing between the $E_{1}$ and $H_{1}$ subbands at $k_{z}\!=\!0$ is essentially caused by the anisotropic Hamiltonian $H^{\prime\prime}$.
@@
We next apply the unitary transformation to the $k_{z}\!\neq\!0$ part of the effective Hamiltoinan $H_{+}(k_{z})$. Combining the result of $k_{z}\!=\!0$ part with that of the $k_{z}\!\neq\!0$ part, we have the transformed effective Hamiltonian
@@
We find that the transformed effective Hamiltonian has the same form as the first term of $H_{+}(k_{z})$ before the transformation [see Eq. (27)]. Recall that the first term of $H_{+}(k_{z})$ is actually the effective Hamiltonian in the isotropic approximation [23]. In this respect, the effect of the anisotropic Hamiltonian $H^{\prime\prime}$ can be regarded as a renormalization to the parameters $A\ldots\,G$ and $\Delta_{e,h_{1,2}}$ of the effective Hamiltonian in the isotropic approximation. The renormalized parameters of the effective Hamiltonian near the gap closing-and-reopening transition are given in Tab. 2.
@@ V The effect of the bulk inversion asymmetry
where $B_{\pm}$ and $\varkappa_{0}$ are material parameters characterizing the strength of the inversion asymmetry. Typical values of these parameters for HgTe can be found in Ref. [37]. However, as we will show in the following, we do not need to know these values here. The bulk inversion asymmetry does not give rise to spin splitting in the [001] oriented cylindrical HgTe nanowire.
@@
Our case is not the first in which the bulk inversion asymmetry does not give rise to spin splitting. In three dimension, the bulk inversion asymmetry does not give rise to spin splitting for both electrons and holes when $\bf k$} is parallel to [100] [28]. In quasi-two dimension, e.g., in a symmetric quantum well grown along the [110] direction, the bulk inversion asymmetry also does not give rise to spin splitting when $\bf k$$\parallel$[001] [28, 39]. In quasi-one dimension, we speculate the cylindrical shape of the nanowire, i.e., its circular cross-section, plays an important role for the vanishing of the spin splitting. When the nanowire has a rectangular cross-section, as studied in [48], the bulk inversion asymmetry does give rise to spin splitting.
VI Summary
The impacts of both the anisotropic term and the bulk inversion asymmetry term of the Kane model on the subband structure of a cylindrical HgTe nanowire are studied in detail. Due to the cylindrical geometry of the nanowire, the bulk inversion asymmetry does not give rise to spin splitting in the subbands. The anisotropic term leads to an anticrossing between the $E_{1}$ and $H_{1}$ subbands at $k_{z}\!=\!0$, and the position of the gap-closing-and-reopening transition has moved to a $k_{z}\!\neq\!0$ place. The low-energy effective Hamiltonian is a six-band model with two separate $3\!\times\!3$ blocks. The [001] oriented HgTe nanowire is in the topological insulator regime when its radius is larger than 3.45 nm.

### v2

@@ IV Transformation of representation
We find $\Delta^{\prime}_{e}$ is always larger than $\Delta^{\prime}_{h1}$, regardless of the detailed values of $\Delta_{e,h_{1}}$ and $\zeta_{e,h_{1},eh_{1}}$. This result has actually been manifested in Fig. 1(b), where the $E_{1}$ energies always lie above the $H_{1}$ energies at $k_{z}\!=\!0$ when we tune the nanowire radius $R$. Hence, the anticrossing between the $E_{1}$ and $H_{1}$ subbands at $k_{z}\!=\!0$ is essentially caused by the anisotropic Hamiltonian $H^{\prime\prime}$.
@@
We next apply the unitary transformation to the $k_{z}\!\neq\!0$ part of the effective Hamiltonian $H_{+}(k_{z})$. Combining the result of $k_{z}\!=\!0$ with that of the $k_{z}\!\neq\!0$, we have the following transformed effective Hamiltonian
@@
We find that the transformed effective Hamiltonian has the same form as the first term of $H_{+}(k_{z})$ before the transformation [see Eq. (27)]. Recall that the first term of $H_{+}(k_{z})$ is actually the effective Hamiltonian in the isotropic approximation [24]. In this respect, the effect of the anisotropic Hamiltonian $H^{\prime\prime}$ can be regarded as a renormalization to the parameters $A\ldots\,G$ and $\Delta_{e,h_{1,2}}$ of the effective Hamiltonian in the isotropic approximation. The renormalized parameters of the effective Hamiltonian near the gap closing-and-reopening transition are given in Tab. 2. Note that when the nanowire radius increases, the sign of parameter AA’ changes (from positive to negative) at about R=3.35R\!=\!3.35 nm, while the sign of parameter $B$’ changes (from negative to positive) at about R=3.50R\!=\!3.50 nm. The sign changes of AA’ and $B$’ do not occur simultaneously. This is possibly caused by the nonzero wave vector position ($k_{z}R\!\neq\!0$) of the gap-closing-and-reopening transition.
@@ V The effect of the bulk inversion asymmetry
where $B_{\pm}$ and $\varkappa_{0}$ are material parameters characterizing the strength of the inversion asymmetry and h.c. stands for Hermitian conjugation. Typical values of these parameters for HgTe can be found in Ref. [38]. However, as we will show in the following, we do not need to know these values here. The bulk inversion asymmetry does not give rise to spin splitting in the [001] oriented cylindrical HgTe nanowire.
@@
Our case is not the first in which the bulk inversion asymmetry does not give rise to spin splitting. In three dimension, the bulk inversion asymmetry does not give rise to spin splitting for both electrons and holes when $\bf k$} is parallel to [001] [29]. In quasi-two dimension, e.g., in a symmetric quantum well grown along the [110] direction, the bulk inversion asymmetry also does not give rise to spin splitting when $\bf k$$\parallel$[001] [29, 40]. The vanishing of the spin splitting is due to the $C_{2v}$ symmetry group of the wave vector $\bf k$} that parallel to [001] [29]. This group has only one irreducible double-group representation $\Gamma_{5}$ which is two-dimensional [29]. For quasi-one dimensional HgTe nanowires grown along the [001] direction, both the square and the circular cross-sections preserve the $C_{2v}$ symmetry, such that in our case the spin splitting vanishes. While the rectangular cross-section breaks the $C_{2v}$ symmetry to $C_{2}$ symmetry, as studied in Ref. [49], the bulk inversion asymmetry leads to spin splitting.
VI Discussion and summary
Here, we briefly discuss the potential influences of the higher-order terms of $k_{z}$ and $\gamma_{3}\!-\!\gamma_{2}$ on the band inversion transition. These higher-order terms may change both the gap-closing-and-reopening position and the anticossing size in the subband structure, but we believe the qualitative behavior of the gap-closing-and-reopening transition will not be influenced. The occurrence of the gap-closing-and-reopening transition in tuning the radius is a consequence of the negative bulk band gap of HgTe (or $\alpha$-Sn), i.e., the negative $\varepsilon_{0}$ in Eq. (1).
The impacts of both the anisotropic term and the bulk inversion asymmetry term of the Kane model on the subband structure of a [001] oriented cylindrical HgTe nanowire are studied in detail. Due to the cylindrical geometry of the nanowire, the bulk inversion asymmetry does not give rise to spin splitting in the subbands. The anisotropic term leads to an anticrossing between the $E_{1}$ and $H_{1}$ subbands at $k_{z}\!=\!0$, and the position of the gap-closing-and-reopening transition has moved to a $k_{z}\!\neq\!0$ place. The low-energy effective Hamiltonian is a six-band model with two separate $3\!\times\!3$ blocks. The [001] oriented HgTe nanowire is in the topological insulator regime when its radius is larger than 3.45 nm.